Main INDEX
Monthly INDEX
PREV
NEXT
User name LeRose
Log entry time 00:03:50 on April 25, 2008
Entry number 234206
keyword=Shift Summary (swing)
Swing Shift 4/24/08
J. LeRose (SL)
L. Mercado (TO)
T. Jinasundera (3rd)
Plan: Take production data at these settings all night.
Kinematics J:
Left Arm: 16.5 degrees, 0.986 GeV/c
Right Arm: 105 degrees, 0.4596 GeV/c
BigBite: Iset=390 Ireadback=388.1 43.6degrees
Beam 1.192 GeV
LH2 target
Shift starts with Hall in Controlled access for He bag
refilling
see halog 234129 for before and after plots
16:10 Out of access
16:19 beam back at 4.4 microAmps
22:10 ROC5 problems. Rebooted ROC5 and restarted CODA
time run# comments
16:12 3739 Test Run
16:12 3740 Test Run
16:21 3741 Production parity DAQ running during the
1st 140k events
16:25 20749 Production, mislabeled in start run
16:31 3742 Production
16:41 3743 Production
16:51 3744 Production
17:03 3745 Production
17:13 3746 Production
17:24 3747 Production
17:35 3748 Production (looks odd on replay.
5 minutes of no beam during
the 15 minute run.
See halog 234175)
17:50 3749 Production
18:00 3750 Production
18:11 3751 Production
18:21 3752 Production
18:32 3753 Production
18:42 3754 Production
18:44 20750 Production
18:55 3755 Production
19:08 3756 Production
19:18 3757 Production
19:37 3758 Production
19:49 3759 Production
20:00 3760 Production
20:11 3761 Production
20:22 3762 Production
20:32 3763 Production
20:43 3764 Production
30:46 20751 Production
20:54 3765 Production
21:05 3766 Production
21:15 3767 Production
21:30 3768 Production
21:43 3769 Production
21:54 3770 Production (Roc 5 problem)
22:36 3771 Production
22:40 20752 Production
22:47 3772 Production
22:58 3773 Production
23:09 3774 Production
23:21 3775 Production
23:32 3776 Production
23:44 3777 Production
23:54 3778 Production
Pretty smooth running